# Absolutely Maximal Contextual Correlations

### v1

@@ III.1 $(3,2,2)$-Bell scenario
$P(x_{1}x_{2}x_{3}|X_{1}X_{2}X_{3})=\begin{cases}\frac{1}{4},&x_{1}\oplus x_{2}\oplus x_{3}=1\oplus X_{1}X_{2}X_{3}\oplus X_{1}X_{2}\oplus X_{2}X_{3}\oplus X_{3}X_{1},\\[6.0pt] \frac{1}{8},&\text{for }(X_{1},X_{2},X_{3}^{\prime}),\;(X_{1},X_{2}^{\prime},X_{3}),\;(X_{1}^{\prime},X_{2},X_{3}),\;(X_{1}^{\prime},X_{2}^{\prime},X_{3}^{\prime}),\\[6.0pt] 0,&\text{otherwise},\end{cases}$ (20)
@@
$\begin{split}p(x_{1}x_{2}|X_{1}X_{2})&=\frac{1}{4}\left[\delta_{0}^{x_{1}}\delta_{0}^{x_{2}}+\delta_{0}^{x_{1}}\delta_{1}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{0}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{1}^{x_{2}}\right],\\ p(x_{1}x_{3}|X_{1}X_{3})&=\frac{1}{4}\left[\delta_{0}^{x_{1}}\delta_{0}^{x_{3}}+\delta_{0}^{x_{1}}\delta_{1}^{x_{3}}+\delta_{1}^{x_{1}}\delta_{0}^{x_{3}}+\delta_{1}^{x_{1}}\delta_{1}^{x_{3}}\right],\\ p(x_{2}x_{3}|X_{2}X_{3})&=\frac{1}{4}\left[\delta_{0}^{x_{2}}\delta_{0}^{x_{3}}+\delta_{0}^{x_{2}}\delta_{1}^{x_{3}}+\delta_{1}^{x_{2}}\delta_{0}^{x_{3}}+\delta_{1}^{x_{2}}\delta_{1}^{x_{3}}\right],\end{split}$ (21)
@@
$p(x_{1}x_{2}x_{3}|X_{1}X_{2}X_{3})=\begin{cases}1/4,&\text{if }x_{1}\oplus x_{2}\oplus x_{3}=X_{1}X_{2}X_{3},\\ 0,&\text{otherwise.}\end{cases}$ (22)
@@
There exist cases in the $(3,2,2)$-Bell scenario that are maximally contextual but do not qualify as AMCCs, as they fail to satisfy the maximal marginals condition. For instance, consider the correlations presented in Table 1: their bipartite marginals are not uniformly random.
@@
We present additional instances of AMCC for the $(3,2,2)$ scenario here. We provide a general construction and will extend it to broader families in the next section. We begin by considering a parametrized family of no-signaling $(3,2,2)$ Bell scenarios that satisfy the normalization condition. The corresponding polytope has dimension 26 (reflecting 26 independent parameters), as detailed in Appendix (B). After further parameter reduction, the resulting parametric table for AMCCs is shown in Table 2.
@@
Table 2 now contains only $8$ free parameters. These parameters can only take values in $[0,1/4]$ to preserve the non-negativity of the probabilities. The empirical model corresponding to this table has maximal marginals for all parameter choices.
We examine which conditions on these parameters yield a contextual fraction (CF) of $1$, using the optimization method described earlier. When we fix two of the eight parameters to $0.125$ or $0.25$, and set the remaining parameters to 0, the possible assignments lead to exactly three possible CF values: $\mathrm{CF}\in\{0,0.5,1\}$. These are not the only parameter choices that yield a maximally contextual model. Our results show that if one parameter is set to $0.25$, the remaining seven can take any values in $[0,0.125]$, and the CF remains $1$. Moreover, if two parameters are fixed to any values (including zero) within the allowed range, the CF is always $1$. This demonstrates that the eight-parameter empirical model shown in Table 2, with the parameter combinations described above, yields $\mathrm{CF}=1$ and the maximal marginals by construction. Hence, all such models correspond to AMCCs. We generalize these results in the next section using the Constraint Satisfaction Problem (CSP) approach, employing Table 2 to construct the Boolean empirical model.
@@ IV The Construction of AMC and non-AMC Correlations
Since the evidence provided in the previous section indicates that there exist maximally contextual correlations (MCCs) that are not AMCCs (see Table 1), the relation is asymmetric. Moreover, from the parametric Table 2, we have observed several examples of AMCCs by conditioning on the parameters. In this section, we present systematic constructions that generate both AMCC and non-AMCC  [^2] The correlation with the contextual fraction $\mathrm{CF}=1$ but not the maximal marginal models in a unified framework. We also provide proofs for certain properties of the empirical models introduced in the previous section.
@@ IV.2 Parity Check method
implies that $(X1+X2=X3^{\prime}\pmod{2})$. Combining this with the first parity equation yields $(X3=X3^{\prime}\pmod{2})$, which is a contradiction. Specifically, for the assignments that satisfy the first parity equation and, thus, the corresponding Boolean formula, the second parity equation, together with its associated Boolean formula, becomes unsatisfiable. Hence, this leads to the contradiction shown in Eq. (35), which, in turn, results in the unsatisfiability of the corresponding Boolean equation, Eq. (36). Hence, no consistent assignment exists, i.e., $S_{e}=\varnothing$ for the possibilistic model. Therefore, the empirical model is strongly contextual, and maximal marginality is ensured by the symmetric structure of zero constraints. Specifically, since half of the total Boolean statements are selected for each row, assigning zeros to the absent Boolean statements in the parametric table 4 results in only four nonzero elements per row. To satisfy the normalization condition and the no-signaling constraint, it follows that all nonzero parameters within a row must be equal, which ensures that the bipartite marginals are maximal. Consequently, the correlations associated with the parity equations constitute an AMCC.
@@ IV.3 Construction of AMCC and non-AMCC
$0\leq p_{9}<\tfrac{1}{2},\qquad p_{9}<p_{2}<\tfrac{p_{9}}{2}+\tfrac{1}{4},$
$\qquad 0<p_{11}<\min\!\bigl(p_{2},\ \tfrac{1}{2}-p_{2},2p_{2}-p_{9}\bigr).$
This model is a non‑AMCC because it does not satisfy the maximal‑marginals condition of Definition 7.
The same model becomes an AMCC when $p_{2}=p_{11}=1/4$ and $p_{9}=0$. In the parity‑check method, determining strong contextuality reduces to checking the parity equations, which is computationally straightforward. In contrast, the general CSP‑based method requires verifying both the no‑signaling condition and the satisfiability of the Boolean constraints, adding extra complexity. However, the latter approach yields strongly contextual models with asymmetric probability tables, which are not obtainable from the symmetric parity‑based construction.
@@ VI Conclusion
The results presented here can be extended in future work to include the findings of Ref. Aravinda and Srikanth (2013), where the complementarity trade-off between no-signaling and local indeterminacy for quantum correlations in the $(2,2,2)$ scenario is introduced. Their work presents a stronger version of Bell’s theorem and provides a quantitative definition of a bound on randomness. This scope is justified by the fact that the violation of a generalized Bell inequality up to its algebraic bound occurs if and only if the model is strongly contextual Abramsky et al. (2018). The very existence of non-AMCCs, which are maximally contextual but not maximally marginal, provides a departure from the complementarity seen in the bipartite scenario, thereby offering insights into the deeper relationship between nonlocality, randomness, and other nonclassical features in multipartite correlations.
@@ Appendix A Proper faces and AVN
The proper face $F$ Pironio (2005), with dimension $\dim F\leq\dim P-1$, can be defined by imposing zero constraints:
@@
Applying the same procedure for other inputs yields correlations such that, for given inputs $X_{1},X_{2}$ and outputs $x_{1},x_{2}$, the logical formula $x_{1}\oplus x_{2}=f(X_{1},X_{2})$ must hold with certainty, i.e., $p(x_{1}\oplus x_{2}=f(X_{1},X_{2})\mid X_{1},X_{2})=1$. Any probabilities violating this logical formula are forced to zero by the imposed constraints. Hence, the zero constraints directly correspond to the logical formula and lead to an AVN (all-versus-nothing) proof. These zeros define the face of the polytope with maximally reduced dimension.

### v2

@@ III.1 $(3,2,2)$-Bell scenario
$P(x_{1}x_{2}x_{3}\mid X_{1}X_{2}X_{3})=\begin{cases}\frac{1}{4},&x_{1}\oplus x_{2}\oplus x_{3}=1\oplus X_{1}X_{2}X_{3}\oplus X_{1}X_{2}\oplus X_{2}X_{3}\oplus X_{3}X_{1},\\[6.0pt] \frac{1}{8},&\text{for }(X_{1},X_{2},X_{3}^{\prime}),\;(X_{1},X_{2}^{\prime},X_{3}),\;(X_{1}^{\prime},X_{2},X_{3}),\;(X_{1}^{\prime},X_{2}^{\prime},X_{3}^{\prime}),\\[6.0pt] 0,&\text{otherwise},\end{cases}$ (20)
@@
$\begin{split}p(x_{1}x_{2}\mid X_{1}X_{2})&=\frac{1}{4}\left[\delta_{0}^{x_{1}}\delta_{0}^{x_{2}}+\delta_{0}^{x_{1}}\delta_{1}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{0}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{1}^{x_{2}}\right],\\ p(x_{1}x_{3}\mid X_{1}X_{3})&=\frac{1}{4}\left[\delta_{0}^{x_{1}}\delta_{0}^{x_{3}}+\delta_{0}^{x_{1}}\delta_{1}^{x_{3}}+\delta_{1}^{x_{1}}\delta_{0}^{x_{3}}+\delta_{1}^{x_{1}}\delta_{1}^{x_{3}}\right],\\ p(x_{2}x_{3}\mid X_{2}X_{3})&=\frac{1}{4}\left[\delta_{0}^{x_{2}}\delta_{0}^{x_{3}}+\delta_{0}^{x_{2}}\delta_{1}^{x_{3}}+\delta_{1}^{x_{2}}\delta_{0}^{x_{3}}+\delta_{1}^{x_{2}}\delta_{1}^{x_{3}}\right],\end{split}$ (21)
@@
$p(x_{1}x_{2}x_{3}\mid X_{1}X_{2}X_{3})=\begin{cases}1/4,&\text{if }x_{1}\oplus x_{2}\oplus x_{3}=X_{1}X_{2}X_{3},\\ 0,&\text{otherwise.}\end{cases}$ (22)
@@
There exist cases in the $(3,2,2)$-Bell scenario that are maximally contextual but do not qualify as AMCCs, as they fail to satisfy the maximal marginals condition. For instance, consider the correlations presented in Table (1): their bipartite marginals are not uniformly random.
@@
We present additional instances of AMCC for the $(3,2,2)$ scenario here. We provide a general construction and will extend it to broader families in the next section. We begin by considering a parametrized family of no-signaling $(3,2,2)$ Bell scenarios that satisfy the normalization condition. The corresponding polytope has dimension 26 (reflecting 26 independent parameters), as detailed in Appendix (B). After further parameter reduction, the resulting parametric table for AMCCs is shown in Table (2).
@@
Table (2) now contains only $8$ free parameters. These parameters can only take values in $[0,1/4]$ to preserve the non-negativity of the probabilities. The empirical model corresponding to this table has maximal marginals for all parameter choices.
We examine which conditions on these parameters yield a contextual fraction (CF) of $1$, using the optimization method described earlier. When we fix two of the eight parameters to $0.125$ or $0.25$, and set the remaining parameters to 0, the possible assignments lead to exactly three possible CF values: $\mathrm{CF}\in\{0,0.5,1\}$. These are not the only parameter choices that yield a maximally contextual model. Our results show that if one parameter is set to $0.25$, the remaining seven can take any values in $[0,0.125]$, and the CF remains $1$. Moreover, if two parameters are fixed to any values (including zero) within the allowed range, the CF is always $1$. This demonstrates that the eight-parameter empirical model shown in Table (2), with the parameter combinations described above, yields $\mathrm{CF}=1$ and the maximal marginals by construction. Hence, all such models correspond to AMCCs. We generalize these results in the next section using the Constraint Satisfaction Problem (CSP) approach, employing Table (2) to construct the Boolean empirical model.
@@ IV The Construction of AMC and non-AMC Correlations
Since the evidence provided in the previous section indicates that there exist maximally contextual correlations (MCCs) that are not AMCCs (see Table (1)), the relation is asymmetric. Moreover, from the parametric Table (2), we have observed several examples of AMCCs by conditioning on the parameters. In this section, we present systematic constructions that generate both AMCC and non-AMCC  [^2] The correlation with the contextual fraction $\mathrm{CF}=1$ but not the maximal marginal models in a unified framework. We also provide proofs for certain properties of the empirical models introduced in the previous section.
@@ IV.2 Parity Check method
implies that $(X1+X2=X3^{\prime}\pmod{2})$. Combining this with the first parity equation yields $(X3=X3^{\prime}\pmod{2})$, which is a contradiction. Specifically, for the assignments that satisfy the first parity equation and, thus, the corresponding Boolean formula, the second parity equation, together with its associated Boolean formula, becomes unsatisfiable. Hence, this leads to the contradiction shown in Eq. (35), which, in turn, results in the unsatisfiability of the corresponding Boolean equation, Eq. (36). Hence, no consistent assignment exists, i.e., $S_{e}=\varnothing$ for the possibilistic model. Therefore, the empirical model is strongly contextual, and maximal marginality is ensured by the symmetric structure of zero constraints. Specifically, since half of the total Boolean statements are selected for each row, assigning zeros to the absent Boolean statements in the parametric table (4) results in only four nonzero elements per row. To satisfy the normalization condition and the no-signaling constraint, it follows that all nonzero parameters within a row must be equal, which ensures that the bipartite marginals are maximal. Consequently, the correlations associated with the parity equations constitute an AMCC.
@@ IV.3 Construction of AMCC and non-AMCC
$0\leq p_{2}<\tfrac{1}{2},\qquad p_{2}<p_{1}<\tfrac{p_{2}}{2}+\tfrac{1}{4},$
$\qquad 0<p_{3}<\min\!\bigl(p_{1},\ \tfrac{1}{2}-p_{1},2p_{1}-p_{2}\bigr).$
This model is a non‑AMCC because it does not satisfy the maximal‑marginals condition of Definition (7).
The same model becomes an AMCC when $p_{1}=p_{3}=1/4$ and $p_{2}=0$. In the parity‑check method, determining strong contextuality reduces to checking the parity equations, which is computationally straightforward. In contrast, the general CSP‑based method requires verifying both the no‑signaling condition and the satisfiability of the Boolean constraints, adding extra complexity. However, the latter approach yields strongly contextual models with asymmetric probability tables, which are not obtainable from the symmetric parity‑based construction.
@@ VI Conclusion
The results presented here can be extended in future work to include the findings of Ref. S and Radhakrishna (2015), where the complementarity trade-off between no-signaling and local indeterminacy for quantum correlations in the $(2,2,2)$ scenario is introduced. Their work presents a stronger version of Bell’s theorem and provides a quantitative definition of a bound on randomness. This scope is justified by the fact that the violation of a generalized Bell inequality up to its algebraic bound occurs if and only if the model is strongly contextual Abramsky et al. (2018). The very existence of non-AMCCs, which are maximally contextual but not maximally marginal, provides a departure from the complementarity seen in the bipartite scenario, thereby offering insights into the deeper relationship between nonlocality, randomness, and other nonclassical features in multipartite correlations.
@@ Appendix A Proper faces and AVN
The proper face $F$ Pironio (2005); Rai et al. (2019), with dimension $\dim F\leq\dim P-1$, can be defined by imposing zero constraints:
@@
Applying the same procedure for other inputs yields correlations such that, for given inputs $X_{1},X_{2}$ and outputs $x_{1},x_{2}$, the logical formula $x_{1}\oplus x_{2}=f(X_{1},X_{2})$ must hold with certainty, i.e., $p(x_{1}\oplus x_{2}=f(X_{1},X_{2})\mid X_{1},X_{2})=1$. Any probabilities violating this logical formula are forced to zero by the imposed constraints. Hence, the zero constraints directly correspond to the logical formula and lead to an AVN (all-versus-nothing) proof Liu et al. (2024). These zeros define the face of the polytope with maximally reduced dimension.